# Entity-aware Cross-lingual Claim Detection for Automated Fact-checking

### v1

@@ Introduction
We report the accuracy, precision, recall, and F1-score of the claim detection models. We fine-tuned ten models and report the average and standard deviations for robust comparison. Since each language is composed of different ratios of verifiable and unverifiable claims, we computed the weighted precision, recall, and F1-score at the language level, and the overall performance is obtained by averaging the language-level performance of the models.
Baseline models
@@
In the absence of cross-lingual claim detection models in the literature, we leverage strong cross-lingual language models as our baselines. We compare mBERT (Alam et al. 2021b; Uyangodage, Ranasinghe, and Hettiarachchi 2021; Panda and Levitan 2021; Zengin, Kartal, and Kutlu 2021; Hasanain and Elsayed 2022), XLM-RoBERTa (XLM-R) (Alam et al. 2021b, a), and mT5 (Du, Gollapalli, and Ng 2022) fine-tuned on the CheckThat! Train partition as the baselines. These models are widely used as a solution to handle multilingual claims and have shown to be consistently competitive across claim detection studies. In addition to these transformer-based language models, we compare some of the open-sourced large language models experimented in Li, Panchendrarajan, and Zubiaga (2024) as the baselines. This includes Llama2-7B, Mistral-7B, and Phi3 fine-tuned on the same training data for the text generation task. Refer to Appendix C for more details on the prompt used and the LLM fine-tuning process.
Results
Baseline Entity Type Categories
Language
mBERT
XLM-R
mT5
Llama2
Mistral
Phi3
X-Claim
EX-Claim-C
EX-Claim-CP
EX-Claim-F
EX-Claim-FP
Arabic 0.692 0.654 0.677 0.752 0.736 0.718 0.731 0.733 0.736 0.733 0.735
Bengali 0.725 0.655 0.556 0.697 0.642 0.552 0.713 0.759 0.755 0.748 0.764
Bulgarian 0.783 0.633 0.729 0.835 0.815 0.812 0.782 0.787 0.789 0.79 0.796
Czech 0.707 0.647 0.72 0.808 0.786 0.755 0.741 0.778 0.775 0.774 0.769
Dutch 0.68 0.698 0.634 0.72 0.657 0.663 0.716 0.721 0.72 0.717 0.718
English (Kaz) 0.738 0.769 0.748 0.764 0.75 0.719 0.755 0.747 0.743 0.757 0.755
English (CT) 0.724 0.658 0.688 0.769 0.777 0.755 0.75 0.758 0.762 0.762 0.761
French 0.638 0.698 0.658 0.736 0.739 0.738 0.703 0.724 0.724 0.722 0.725
German 0.751 0.714 0.722 0.808 0.822 0.816 0.789 0.77 0.767 0.778 0.775
Greek 0.677 0.64 0.71 0.697 0.687 0.679 0.735 0.748 0.746 0.747 0.759
Gujarati 0.661 0.669 0.649 0.457 0.582 0.455 0.771 0.775 0.763 0.77 0.751
Hindi 0.631 0.695 0.613 0.792 0.809 0.6 0.766 0.761 0.751 0.767 0.769
Hungarian 0.667 0.651 0.665 0.757 0.727 0.637 0.68 0.682 0.674 0.683 0.682
Italian 0.721 0.647 0.704 0.764 0.745 0.736 0.709 0.724 0.72 0.731 0.723
Japanese 0.716 0.658 0.693 0.759 0.771 0.74 0.715 0.728 0.728 0.732 0.746
Kannada 0.535 0.567 0.579 0.371 0.459 0.342 0.628 0.666 0.674 0.674 0.676
Korean 0.691 0.635 0.623 0.69 0.72 0.653 0.712 0.724 0.724 0.718 0.709
Marathi 0.588 0.64 0.603 0.643 0.532 0.46 0.668 0.664 0.657 0.666 0.664
Malayalam 0.79 0.564 0.685 0.374 0.151 0.141 0.845 0.826 0.832 0.84 0.835
Persian 0.673 0.669 0.622 0.713 0.667 0.602 0.704 0.718 0.717 0.716 0.718
Russian 0.709 0.667 0.641 0.766 0.764 0.734 0.715 0.721 0.718 0.724 0.724
Spanish 0.738 0.693 0.683 0.779 0.735 0.77 0.787 0.783 0.78 0.783 0.784
Swedish 0.707 0.688 0.654 0.772 0.779 0.76 0.777 0.775 0.776 0.778 0.787
Tamil 0.623 0.744 0.789 0.439 0.398 0.264 0.864 0.853 0.86 0.866 0.861
Thai 0.628 0.648 0.644 0.716 0.698 0.673 0.726 0.711 0.709 0.716 0.715
Turkish 0.699 0.648 0.664 0.775 0.82 0.745 0.748 0.751 0.749 0.759 0.758
Ukrainian 0.688 0.597 0.604 0.764 0.724 0.719 0.644 0.692 0.691 0.688 0.707
Vietnamese 0.727 0.672 0.676 0.735 0.694 0.689 0.715 0.714 0.736 0.717 0.729
Table 5: F1-Score of the Claim Detection Models at Language-level
Dataset-level Performance
Table 4 presents the performance of the claim detection models. The results demonstrate that EX-Claim significantly outperforms the baseline models across all four metrics in overall evaluation. This underscores the importance of learning task-specific representations and attending to critical features to identify the verifiability of a claim. Furthermore, the consistent performance across diverse datasets from various platforms covering multiple topics suggests that our hypothesis is generalizable, regardless of the nature of the claim. It can be observed that the large language models, Llama2 and Mistral achieve the highest performance in CheckThat! dataset. However, their performance drops significantly on the other two datasets, indicating challenges in handling a diverse range of languages. Among the transformer-based baselines, mBERT model outperforms the other two across all datasets. However, utilizing mBERT as the cross-lingual embedding layer in EX-Claim did not yield the optimal performance. This shows that mBERT is powerful in representing cross-lingual context vectors, whereas, XLM-R is good at generating cross-lingual word representations.
Moreover, incorporating entity-type information always enhances the accuracy, precision, recall, and F1-score in CheckThat! and synthetic datasets. This increment is significant in the synthetic dataset composed of 18 unseen languages with at least a 1.2% increment across all the metrics. This shows that, while knowing the context, the EX-Claim utilizes entity information to improve the performance of both seen and unseen languages. Among these, the EX-Claim-FP model which uses fine-grained NER information and EL stands out as the best model across all three datasets. It is interesting to note that the NER models (EX-Claim-C or EX-Claim-F) and their corresponding models utilizing EL information (EX-Claim-FP, EX-Claim-CP) reach very similar performances. This similarity suggests that most of the entities identified as NER are already recognized as popular, resulting in only minor performance improvements.
Language-level Performance
Table 5 lists the language-level F1-score of the claim detection models across 27 languages present in the test datasets. It is noteworthy that large language models (LLMs) such as Llama2 and Mistral achieve the highest performance across high-resource languages. However, their performance declines sharply in various low-resource languages. On the contrary, EX-Claim not only outperforms LLMs in several low-resource languages but also delivers competitive performance in high-resource languages. This highlights that transformer-based models like XLM-R remain highly competitive for text classification tasks, despite their significantly smaller size compared to larger models (e.g., 250M versus 7B parameters). Furthermore, consistent with the overall trends discussed earlier, at least one (although generally all) of the variants of the EX-Claim significantly improves the transformer-based baselines across all the languages except in the English language in the Kazemi dataset. Further, the models that leverage entity information show notable improvements in several languages, with particularly significant gains observed in Bengali, Czech, French, Greek, Italian, Japanese, Kannada, and Ukrainian. However, it is noteworthy that the inclusion of entity information results in a performance decline in certain languages, such as German, Malayalam, and Thai.
Analysis
Cross-lingual Knowledge Transfer
@@ Results
Model
Correct
Prediction
Wrong
Prediction
@@
Entity Types Categories X-Claim 87.21% 71.73%
EX-Claim-C 88.15% 68.76%
EX-Claim-CP 87.43% 68.63%
EX-Claim-F 87.82% 68.48%
EX-Claim-FP 87.09% 66.51%
@@
Table 6 lists the knowledge transfer rate of claim detection models. Here, the transformer-based baseline models exhibit the lowest transfer rate in correct and wrong predictions. In contrast, EX-Claim significantly enhances the transfer rate among the correct and wrong predictions and outperforms the LLMs compared when transferring knowledge across languages in correct prediction. Further, the incorporation of entity information tends to either maintain or improve the transfer rate of correct predictions, while consistently reducing the transfer rate of incorrect predictions. This behavior indicates that the projection layer, which maps input features into a task-specific lower-dimensional vector, effectively transfers knowledge between languages. Additionally, the entity information enhances language-level performance, thereby reducing the transfer rate of incorrect predictions.
Attending to Important Words
@@
EX-Claim-C 3.579
EX-Claim-CP 3.479
EX-Claim-F 3.519
EX-Claim-FP 3.322
@@
Table 7 lists the attention layer’s entropy of the claim detection models. We can see that the models utilizing entity information always yield lower entropy. This shows that these models are concentrated on a few keywords in the input sequence.222 Refer to Appendix B for attention weights generated for sample sentences. Moreover, the entity types identified via the popularity further reduce the entropy of the models.
Entity-type Embedding
(a) EX-Claim-CP
(b) EX-Claim-F
Figure 2: Two-dimensional embedding visualization of EX-Claim-CP and EX-Claim-F Models
Unlike the cross-lingual word embedding which is directly obtained from the pre-trained models, EX-Claim learns a dense vector representation for each entity type. Figures 2(a) and 2(b) show the 2D visualization of the entity type embeddings learned by the EX-Claim-CP and EX-Claim-F models. Popular entities are indicated with the post-fix -P. The 2D vector representation of the embeddings is obtained using the t-SNE [^3] https://scikit-learn.org/stable/modules/generated/sklearn
.manifold.TSNE.html
dimension reduction technique. Further, the embedding representations were clustered using the KMeans clustering technique (Sinaga and Yang 2020), and the number of optimal clusters was chosen using Silhouette score (Rousseeuw 1987).
The EX-Claim-CP model learns three groups of entity types: the location entities form one cluster, both popular and unpopular miscellaneous entities form another cluster, and the remaining entities, irrespective of their popularity, constitute the third cluster. Similarly, EX-Claim-F model learns various groups of entity types. Notably, this model clusters Animal, Disease, Person, Biological Entity, Disease, Plan and Mythological Entity into a single group, likely influenced by the COVID-19 discussions present in the training data. This observation highlights the model’s tendency to learn context-specific entity-type embeddings during training.
Ablation Study
Model F1-Score
X-Claim 0.735
X-Claim - NPL 0.688
X-Claim - NAL 0.714
Table 8: Ablation Results of X-Claim Model
Model F1-Score
EX-Claim-C 0.743
EX-Claim-C - NEL 0.732
EX-Claim-CP 0.742
EX-Claim-CP - NEL 0.731
EX-Claim-F 0.745
EX-Claim-F - NEL 0.733
EX-Claim-FP 0.746
EX-Claim-FP - NEL 0.733
Table 9: Ablation Results of EX-Claim Models
Error Analysis
@@
We analyze the false positive and false negative predictions made by all four variants of EX-Claim. Table 10 lists some of the misclassified examples. We noticed that the appearance of disease names ($S4$), lengthy claims, nonrecognition of personal experience from pronouns’ presence ($S4$), and ambiguous fact expressions could cause false positive scenarios ($S5$).
We observed that misclassification of popular events expressing personal experience tends to be a common cause of false negatives. Refer to Sentences S6-S9 in Table 10 for sample false negative instances where the reach of the event makes it verifiable, yet the personal experience framing leads to misclassification. Further, some of these claims refer to other modalities such as images and videos in the expression of facts, making it further difficult when relying solely on textual information  [^4] Refer to Appendix B showing the correlation between the confusion matrix and some of these features..
Conclusion
We present EX-Claim, the first cross-lingual claim detection model that encapsulates entity information, including their popularity, for improved performance in detecting claims spreading on social media, across multiple languages. Extensive experiments on three datasets comprising 27 languages reveal that EX-Claim significantly outperforms competitive baseline methods. Especially EX-Claim shows competitive performance across high-resource languages and performs better in low-resource languages than the large language models compared. Further, the language-level analysis shows that certain languages particularly benefit from entity information, resulting in improved performance. Furthermore, the models achieve the highest cross-lingual knowledge transfer rate by accurately transferring knowledge from correct predictions while minimizing errors during transfer, a critical capability for addressing misinformation on diverse social media platforms. Our in-depth analysis reveals that EX-Claim achieves the lowest average entropy in attention weights, indicating a focus on key terms to assess verifiability. Future work will explore techniques to generalize entity-type embeddings beyond the context encountered in the training data and address the challenges contributing to false positive and false negative outcomes identified in our error analysis.
Limitations
@@ Appendix A Statistics
Statistics of Named Entity Types Identified
Table 11 lists the types of named entities identified using the coarse-grained and fine-grained NER models. As discussed earlier, WikiNEuRal identifies 4 types of coarse-grained named entities and MultiNERD identifies 15 types of fine-grained Named Entities.
NER Model Entities
WikiNEuRal
Person (PER)
Organization (ORG)
Location (LOC)
Miscellaneous (MISC)
MultiNERD
@@
Table 11: Named Entity types Identified by the NER Models
Figure 3: Average Number of Fine-grained Entities Identified in the Train Partition of CheckThat! 2022 Data
Figure 4: Average Number of Coarse-grained Entities Identified in the Train Partition of CheckThat! Data
Statistics of Named Entities Recognized
Figures 4 and 3 show the average number of coarse-grained and fine-grained named entities recognized respectively in the Train partition of CheckThat! 2022 data. It can be observed that the average number of coarse-grained location entities present in a sentence is relatively higher for all languages compared to the other three types. This could be due to the COVID-19 discussion in the training data referring to the status of the disease from different countries and locations. Miscellaneous entities stand out as the 2nd most frequent coarse-grained entity type for all languages except for Arabic. When comparing coarse-grained and fine-grained NERs, the distribution across languages is similar for the organization entity type. However, both tools recognize different distributions for person and location entity types. Since the dataset contains tweets about COVID-19, the occurrences of disease, media, biological entities, and food are relatively higher compared to other rare entity types.
Statistics of Entities Linked
(a) Coarse-grained EL
(b) Fine-grained EL
@@
(a) Popular vs Unpopular Coarse-grained Entities
(b) Popular vs Unpopular Fine-grained Entities
@@
Figure 5 shows the number of named entities linked to a Wikipedia page with the increase in threshold. It can be observed that the reduction in the number of entities linked to a Wikipedia page with the increase in log probability threshold is not very significant until the threshold is increased to -0.15 (probability score of 0.861). This indicates that most of the coarse-grained or fine-grained NER identified are linked to a Wikipedia page with a very high probability.
Figure 6 shows the fraction of popular and unpopular entities determined with the log probability threshold of -0.15 (probability of 0.861). Both diagram shows a similar pattern, where Arabic language data contains more fraction of entities identified as popular followed by English, and Dutch and Turkish have the lowest fraction among all. Further, the chances of fine-grained entities being popular entities are slightly higher (ranging from 0.39 to 0.63) compared to coarse-grained entities (ranging from 0.33 to 0.61). Moreover, this indicates, that nearly more than 40% of the entities recognized are linked as popular ones in both categories with a very high probability.
Statistics of the Datasets Used
We use two datasets that are publicly available for the experiments, and the datasets may contain personally identifiable information or offensive content.
CheckThat! Dataset

### v3

@@ 4.2 Evaluation metrics
We report accuracy, precision, recall, and F1-score for ten fine-tuned claim detection models, presenting their averages and standard deviations. To account for varying ratios of verifiable and unverifiable claims across languages, we calculated weighted scores at the language level, and the overall performance is obtained by averaging the language-level performance of the models.
4.3 Baseline models
@@
Due to the lack of cross-lingual claim detection models, we use strong cross-lingual language models as baselines: mBERT (Alam et al., 2021b; Uyangodage et al., 2021; Panda and Levitan, 2021; Zengin et al., 2021; Hasanain and Elsayed, 2022), XLM-RoBERTa (XLM-R) (Alam et al., 2021b, a), and mT5 (Du et al., 2022), all fine-tuned on the CheckThat! Train partition. These widely used models have proven competitive in multilingual claim detection. We also include open-source large language models from Li et al. (2024)—Llama2-7B, Mistral-7B, and Phi3—fine-tuned on the same data for text generation.
Refer to Appendix D for resources used for the training, hyperparameters, the prompt used for LLMs, and the LLM fine-tuning process.
5 Results
5.1 Dataset-level Performance
Table 2 summarizes the performance of the claim detection models. Llama2 performs strongly on the CheckThat! 2022 Test dataset, while EXP-Claim achieves the highest accuracy, recall, and F1-score on both the synthetic and Kazemi datasets. The performance of large language models declines on these two datasets, highlighting challenges in handling linguistically diverse data.
Notably, the EX-Claim models and their ablated variants outperform both transformer-based baselines and large language models in overall evaluation, underscoring the value of learning task-specific representations that focus on claim-relevant words. Moreover, their consistent performance across diverse datasets—spanning multiple platforms and topics—suggests that our hypothesis holds generally, regardless of the claim’s nature.
Incorporating NER information (EXN-Claim vs. X-Claim) consistently improves performance across all four metrics. While the overall gain is modest (approximately 1%), our language-level analysis (Section 5.2) shows that NER integration either maintains or enhances model performance across most of the languages, with certain languages benefiting significantly from entity information and linguistic knowledge transfer from the NER and EL tools. Interestingly, incorporating entity popularity (EXN-Claim vs. EXP-Claim) did not lead to further performance improvement—likely because most entities identified by the NER model were already recognized as popular, limiting the potential for additional gains.
5.2 Language-level Gain
(a) Transformer-based Baselines
(b) X-Claim
(c) Large Language Models (LLMs)
Figure 2: Performance Gain in EXN-Claim compared to Three groups of Models
We analyze language-level performance gains from incorporating entity information into claim detection. Since there is no performance gain in EXP-Claim compared to EXN-Claim, we analyze the performance of EXN-Claim with three groups of models as shown in Figure 2. Refer to Appendix B.1 for the language-level F1-score of the models.
Figure 2(a) illustrates the performance gain of the EXN-Claim model over transformer-based baselines. Most languages exhibit consistent positive gains, except English (Kazemi dataset) and Vietnamese, which show slight drops (0.95–1.2%). The most substantial improvements are seen in Asian languages like Malayalam, Tamil, Hindi, Bengali, Kannada, and Gujarati. These results underscore the effectiveness of incorporating entity information and dynamically attending to claim-relevant words, leading to significant performance gains across a diverse set of languages.
Figure 2(b) shows the performance gain of the EXN-Claim model over its ablation variant, X-Claim, which excludes entity information. Interestingly, incorporating entity information either retains or improves model performance across most languages. The only exceptions are German and Thai, which exhibit a slight performance drop of approximately 0.97% to 1.1%. Notably, certain languages—such as Kannada, Ukrainian, Bengali, Czech, Italian, and French—experience the highest gains, ranging from 1.95% to 4.6%. These improvements can likely be attributed to the incorporation of entity information and the indirect linguistic knowledge transfer from the NER tool.
Similarly, we compare the performance gain of the EXN-Claim model with large language models (LLMs) in Figure 2(c). While LLMs outperform EXN-Claim in several high-resource languages, they perform notably poorly in certain low-resource languages—particularly Asian regional languages such as Malayalam, Kannada, Tamil, Gujarati, Bengali, and Hindi. The largest gains are in Malayalam (up to 70%) and Tamil (up to 60%), which include non-synthetic samples from the Kazemi dataset. Despite being smaller and more computationally efficient, EXN-Claim outperforms all three LLMs in 7 languages and surpasses at least one LLM in 16 out of the 27 languages evaluated.
Furthermore, we observe inconsistent gains across LLMs: Llama excels in Dutch, Phi3 in Hungarian, while Mistral underperforms in both. Similarly, Llama and Phi3 perform well in Korean, but Mistral struggles. These observations suggest that LLMs tend to excel only in specific languages, largely influenced by the composition of their training data. In contrast, these findings highlight EXN-Claim as a highly efficient model that consistently delivers superior performance across many low-resource languages, while remaining competitive with LLMs in high-resource scenarios, offering the best balance between overall performance gains and computational efficiency. Refer to Appendix D.3 for comparison of claim detection models based on the number of parameters, multilingual support, and training time.
6 Analysis
6.1 Cross-lingual Knowledge Transfer
@@
Model Correct Prediction Wrong Prediction
@@
X-Claim 87.21% 71.73%
EXN-Claim 87.82% 68.48%
EXP-Claim 87.09% 66.51%
@@
Table 3 lists the knowledge transfer rate of claim detection models. Overall, the transformer-based baseline models exhibit the lowest transfer rate in correct and wrong predictions. In contrast, EX-Claim and its ablated variant achieve the highest transfer rate—reaching nearly 87%—for correct predictions. This highlights the effectiveness of learning context-aware claim representations with attention to claim-relevant words in promoting cross-lingual generalization. Furthermore, the inclusion of entity information through NER and EL (EXN-Claim and EXP-Claim) reduces the transfer of incorrect predictions by 3% to 5%. This indicates that entity-level knowledge and linguistic signals from NER and EL help prevent consistent misclassification of the same instances across different languages, thus improving model robustness.
6.2 Attending to Important Words
@@
EX-Claim 3.519
EXP-Claim 3.322
@@
Table 4 lists the attention layer’s entropy of the claim detection models. We can see that the models utilizing entity information always yield lower entropy, especially when utilizing popularity, resulting in the lowest entropy among all. This shows that these models are concentrated on a few keywords in the input sequence to determine the verifiability of a claim. [^2] Refer to Appendix C.2 for attention weights generated for sample sentences.
6.3 Error Analysis
@@
We analyze the false positive and false negative predictions made by EX-Claim models. Table 5 lists some of the misclassified examples. We noticed that the appearance of disease names ($S4$), lengthy claims, nonrecognition of personal experience from pronouns’ presence ($S4$), and ambiguous fact expressions could cause false positive scenarios ($S5$).
We observed that misclassification of popular events expressing personal experience tends to be a common cause of false negatives. Refer to Sentences S6-S9 in Table 5 for sample false negative instances where the reach of the event makes it verifiable, yet the personal experience framing leads to misclassification. Further, some of these claims refer to other modalities such as images and videos in the expression of facts, making it further difficult when relying solely on textual information  [^3] Refer to Appendix C.4 showing the correlation between the confusion matrix and some of these features..
7 Conclusion
We present EX-Claim, the first cross-lingual claim detection model that integrates entity information for improved performance in detecting claims across multiple languages. Extensive experiments on three datasets comprising 27 languages reveal that EX-Claim stands out as the effective model with consistent performance gain across multiple languages compared to the baselines. Language-level performance analysis reveals that certain languages—particularly Asian regional languages—benefit significantly from the incorporation of entity information. This improvement is likely driven by the emphasis on key terms essential for claim detection and by the indirect transfer of linguistic knowledge from named entity recognition and entity linking tools. Further analysis of cross-lingual transfer rates and attention layer entropy supports this finding. Notably, EX-Claim achieves the highest cross-lingual knowledge transfer rate across correct predictions while minimizing the wrong predictions across claims written in the same language, demonstrating its robustness in cross-lingual generalization and transfer efficiency. Future work will focus on generalizing entity-type embeddings beyond training contexts and address the challenges contributing to false-positive and false-negative outcomes identified in our error analysis.
8 Limitations
@@ Appendix A Statistics
Figure 3: Average Number of Named Entities Identified in the Train Partition of CheckThat! 2022 Data
A.1 Statistics of Named Entities Recognized
MultiNERD identifies the following 15 types of fine-grained Named Entities.
@@ A.1 Statistics of Named Entities Recognized
Figures 3 show the average number of named entities recognized in the Train partition of CheckThat! 2022 data. Notably, location entities appear more frequently across all languages. This could be due to the COVID-19 discussion in the training data referring to the status of the disease from different countries and locations. Additionally, the occurrences of disease, media, biological entities, and food are relatively higher compared to other rare entity types, likely due to the COVID-19-related discussions present in the dataset.
A.2 Statistics of Entities Linked
@@
Figure 5: Popular vs Unpopular Fine-grained Entities
@@
Figure 4 shows the number of named entities linked to a Wikipedia page with the increase in threshold. It can be observed that the reduction in the number of entities linked to a Wikipedia page with the increase in log probability threshold is not very significant until the threshold is increased to -0.15 (probability score of 0.861). This indicates that most of the NER identified are linked to a Wikipedia page with a very high probability.
Figure 6 shows the fraction of popular and unpopular entities determined with the log probability threshold of -0.15 (probability of 0.861). Notably, Arabic language data contains more fraction of entities identified as popular followed by English, and Dutch and Turkish have the lowest fraction among all. Further, the chances of fine-grained entities being popular range from 0.39 to 0.63, highlighting that nearly more than 40% of the entities recognized are linked as popular ones with a very high probability.
A.3 Statistics of the Datasets Used
A.3.1 CheckThat! Dataset